How much time does a meter stick moving at 0.1c relative to an observer take to pass the


observer? The meter stick is parallel to its direction of motion.

1
Expert's answer
2022-04-27T13:34:32-0400
"l=l_0\sqrt{1-(v/c)^2}"

"l=\rm 1.00\: m\times\sqrt{1-(0.1)^2}=0.995\:m"